DISPERSED FLUORESCENCE SPECTROSCOPY OF THE B̃ ^{2}E′ – X̃ ^{2}A′_{2} TRANSITION OF NO_{3}

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated NO_{3} in supersonic free jet expansions and observed laser induced fluorescence ( LIF ) of the B̃ ^{2}E′ – X̃ ^{2}A′_{2} transition. We have measured LIF excitation spectrum and dispersed fluorescence ( DF ) spectra from the single vibronic levels ( SVL ) of the B̃ ^{2}E′ state. Comparing with the vibrational structures of the DF spectra, vibrational structure of the X̃ ^{2}A′_{2} state has been analyzed. It is especially concluded that the 1,492 cm^{−1} level in the X̃ ^{2}A′_{2} state has no contribution of the ν_{4} ( e′ ) mode, though there are two opinions in which the level is attributed to the ν_{3} ( e′ ) fundamental or the ν_{3} + ν_{4} combination level.